Your uncle offers to take you on an airplane ride. The plane has an open cockpit. Your mother wants you to wear a warm jacket, even though it is summer and the temperature at the airport is 75 degrees. You mother knows that the air temperature drops 2o for every 530 feet you climb in altitude in the plane. If you uncle plans to fly up to 10,600 feet, how cold will it be in the open cockpit?

Answer by checkley77(12569)

75-2(10,600/530)=75-2(20)=75-40=35 degrees @ 10,600 ft.